OpenStudy (anonymous):

Simplify. (will post everything, what to simplify and answer choices inside!) Thanks! :) please explain :)

OpenStudy (anonymous):



OpenStudy (anonymous):

answer choices A,B,C,D from top to bottom :)

OpenStudy (anonymous):

\[-\sqrt{-1}\sqrt{2}\sqrt{100}\]

OpenStudy (anonymous):

\[-i(10)\sqrt{2}\]

OpenStudy (anonymous):